What is the main difference between average physical and marginal physical product? Give an example of marginal physical property.
Answer #1

Average physical product

The quantity of total units produced per unit of variable input holding other input fixed is called Average physical product(APP)

Formula = TPP(Total physical product)/ no of variable input

APP is found by dividing total physical product by number of variable input.

Marginal physical product

The change in quantity of total physical product resulting from unit change in variable input keeping variable inputs unchanged.

MPP is found by using

MPPn+1 = TPPn+1 - TPPn

Suppose

( Refer to table below.)

n = 1

then

MPP2 = TPP2 - TPP1

   = 30-10

= 20

No of variable input TPP APP MPP
1 10 10 (10/1) 10
2 30 15(30/2) 20 (30-10)
3 60 20(60/3) 30(60-30)
4 76 19(76/4) 16(76-60)
5 70 14(70/5) -6 (70-76)

ALL THE BEST :)))